# Spatial Context-Driven Positive Pair Sampling for Enhanced Histopathology Image Classification

### v1

@@ 3.5.1 Proportion of Contextual Similarity
Our loss formulation balances contextual similarity and transformation invariance, weighted by $\alpha$. Figure 3 shows the effect of such proportion on the classification accuracy for all the datasets. The data indicates that incorporating the contextual information ($\alpha>0$) yields accuracy gains relative to a baseline ($\alpha=0$). Moreover, while $\alpha=1$ (only neighbor-based pairs) outperforms $\alpha=0$, it is rarely the absolute best. In most datasets and methods, a mixture of contextual and standard augmentation-based pairs yields the highest accuracy. However, the magnitude and behavior of these gains vary by both method and dataset. For instance, BYOL shows more variable performance, with sensitivity to the $\alpha$ value, while VICReg generally provides consistent improvements. Overall, these results validate the key idea that balancing “transformation invariance” and “contextual similarity” is beneficial. The exact sweet spot for $\alpha$ varies, but typically falls in the middle range, underscoring the importance of simultaneously leveraging both types of positive pairs in self-supervised histopathology image analysis.
Figure 3: Accuracy gain relative to the baseline (without contextual information; i.e., $\alpha=0$) for each SSL method across four datasets. Results are averaged over five runs.
@@ 3.5.2 Effect of Neighboring Distance
We investigated how distance thresholds for selecting contextual positive pairs affect SSL performance in slide-level classification. We sampled 100,000 pivot patches and identified neighbor patches within specific distance thresholds, covering both short- and long-range contexts.
Figure 4 illustrates how neighboring distance affects classification accuracy across all datasets. Overall, shorter distances yield higher accuracy gains, while performance typically declines when the distance exceeds 5—likely due to diminished contextual relevance and increased noise or mismatches. Although all three SSL methods benefit from closer patch sampling, the magnitude of improvement varies with dataset-specific factors such as tissue heterogeneity and the SSL method employed. Interestingly, accuracy in the Stomach and Colon datasets rebounds at greater distances, a trend we attribute to these WSIs containing multiple tissue samples. Based on these observations and our false positive pair analysis, we recommend sampling neighbor patches immediately adjacent to the pivot (i.e., $distance=1$).
@@ 4 Discussion
We introduced a spatial context-driven positive pair sampling strategy that enhances self-supervised learning (SSL) in whole-slide image (WSI) analysis by leveraging the inherent spatial coherence of tissue morphology. Our experimental results show that integrating contextual pairs improves slide-level classification under Multiple Instance Learning, with ablation studies indicating that smaller neighboring distances generally yield greater gains—although the optimal distance varies by dataset. Excessive reliance on spatial proximity alone can reduce representation quality, highlighting the need to balance contextual and standard augmentations. While our study did not aim to surpass state-of-the-art performance benchmarks, we focused on establishing a biologically meaningful SSL framework that improves representation learning in computational histopathology. Our findings underscore the importance of spatial context in guiding SSL models toward more diagnostically relevant feature extraction, which could be particularly beneficial for training Pathology Foundation Models on large-scale, annotation-limited datasets. Additionally, optimizing our approach for computational efficiency and exploring its applicability across diverse tissue types and staining variations remain key areas for future research. In conclusion, by aligning positive pair formation with the inherent structure of WSIs, our method provides a promising direction for self-supervised histopathology analysis, paving the way for more scalable and generalizable AI-driven pathology models.

### v2

@@ 4.4 Slide-level Classification
4.5 Linear Probing Evaluation
Table 2: Patch-level performance (Linear Probing) across datasets, reporting accuracy and AUROC for positive pair sampling strategies. Standard refers to self-supervised learning without contextual pairing (i.e., only augmentation-based positive pairs), while Context($1$) refers to our method using spatially adjacent pairs (Chebyshev distance of 1). Supervised indicates fully supervised training; the first value corresponds to a ResNet-18 backbone, and the second (in the DINOv2 row) to a ViT-Tiny backbone. Results are averaged over five runs.
Dataset SSL Accuracy AUROC
Supervised Standard Context($1$) Supervised Standard Context($1$)
Private (Stomach) BT 0.828 0.739 0.783 0.942 0.912 0.928
BYOL 0.774 0.809 0.913 0.938
VICReg 0.803 0.821 0.933 0.944
DINOv2 0.830 0.750 0.762 0.944 0.896 0.904
Private (Colon) BT 0.818 0.840 0.827 0.943 0.939 0.935
BYOL 0.861 0.885 0.950 0.963
VICReg 0.859 0.862 0.950 0.952
DINOv2 0.878 0.797 0.858 0.962 0.921 0.945
Camelyon16 (Breast) BT 0.885 0.764 0.756 0.947 0.831 0.817
BYOL 0.777 0.782 0.834 0.821
VICReg 0.792 0.798 0.860 0.877
DINOv2 0.910 0.797 0.858 0.964 0.921 0.945
To further evaluate the quality of the learned representations, we conducted linear probing using both ResNet-18 and ViT-Tiny backbones. As shown in Table 2, incorporating spatially contextual positive pairs (Context($1$)) consistently improves performance over standard SSL across most datasets and methods.
In the Stomach and Colon datasets, contextual pairing improves AUROC and accuracy in nearly all cases, especially for BYOL and VICReg, which benefit from the spatial coherence in tissue morphology. Notably, BYOL on Colon improves from 0.861 to 0.885 in accuracy and from 0.950 to 0.963 in AUROC. Similar trends are observed in VICReg and DINOv2. For the Camelyon16 (Breast) dataset, improvements are more modest and less consistent. In some cases, such as BT, contextual sampling slightly reduces performance, suggesting that local morphological context may be less discriminative in this dataset, possibly due to higher heterogeneity in slide content or smaller ROI areas.
Across all datasets, standard SSL representations perform worse than both the supervised baseline and our contextual method, reinforcing that augmentation-based invariance alone may be insufficient for modeling histopathology structures. Importantly, contextual SSL nearly closes the gap with fully supervised training and, in some cases (e.g., BYOL and VICReg on Colon), outperforms the supervised ResNet-18 model.
Finally, ViT-Tiny models (DINOv2 rows) exhibit notable improvements with contextual pairing (e.g., from 0.797 to 0.858 accuracy on Breast), indicating that transformer-based encoders can also benefit from spatial coherence, though they remain more sensitive to context quality.
@@ 4.6.1 Proportion of Contextual Similarity
Our loss balances contextual similarity and transformation invariance through the mixing parameter $\alpha$. Figure 3 shows that incorporating contextual signals ($\alpha>0$) generally improves accuracy across datasets and methods. The best performance typically occurs at intermediate values ($\alpha=0.25$ or $0.5$), while relying solely on contextual pairs ($\alpha=1.0$) often degrades accuracy, particularly for DINOv2. This suggests that contextual cues alone may lack sufficient diversity, especially for transformer-based models.
Among methods, VICReg shows stable improvements, while BYOL is more sensitive to $\alpha$. BT performs well at moderate $\alpha$ values, and DINOv2 benefits from some contextual information but drops sharply when overused.
Figure 3: Accuracy gain relative to the baseline (without contextual information; i.e., $\alpha=0$) for each SSL method across four datasets. Results are averaged over five runs. Dark blue, orange, green, and purple lines correspond to Barlow Twins (BT), BYOL, VICReg, and DINOv2, respectively.
@@ 4.6.2 Effect of Neighboring Distance
We investigated how the spatial distance between pivot and neighbor patches affects the performance of self-supervised learning (SSL) in slide-level classification. From a pool of 100,000 pivot patches, we sampled neighbors within defined distance thresholds to capture both local and broader tissue context. As shown in Figure 4, accuracy gains are highest when neighbor patches are sampled from short distances, with performance peaking consistently at distance 1 across all datasets and SSL methods. Beyond this point, accuracy generally declines, particularly at distances greater than 5, likely due to reduced contextual coherence and increased chances of mismatched or noisy pairs. The effect is consistent across Barlow Twins, BYOL, and VICReg, although the extent of gain varies by dataset and method. This variation may reflect differences in tissue structure and intra-slide heterogeneity.
Interestingly, the Stomach and Colon datasets exhibit a partial rebound in accuracy at the largest tested distance (30), possibly due to the presence of multiple tissue fragments per slide, which can create structurally coherent regions even at wider separations. Nonetheless, our results suggest that immediate spatial proximity (i.e., $distance=1$) offers the most reliable and biologically meaningful context for representation learning.
Due to computational constraints, we excluded DINOv2 from these experiments. Exploring its behavior under spatial proximity-based sampling remains promising future work.
@@ 5 Discussion
5.1 Spatial Context-Driven Positive Pair Sampling
This work presents a spatial context-driven positive pair sampling strategy tailored for self-supervised learning (SSL) in whole-slide image (WSI) analysis. While conventional SSL methods define positive pairs through synthetic augmentations, such approaches often fail to capture the structural continuity and local semantic coherence of histopathology data. Our method addresses this by explicitly incorporating spatial proximity into the positive pair selection process, leveraging the intrinsic grid structure of WSIs to guide representation learning.
Experimental results from both slide-level classification and patch-level linear probing confirm the effectiveness of our approach. Across datasets and SSL frameworks, incorporating contextual positive pairs, especially those sampled from immediate spatial neighbors (Chebyshev distance of 1), consistently improves downstream performance. For instance, Barlow Twins and VICReg show 4 to 8 percent accuracy gains on slide-level classification tasks in the stomach and colon datasets when using spatially adjacent pairs, compared to standard augmentation-only baselines. These improvements suggest that morphologically similar regions in close spatial proximity provide a more informative training signal than synthetic augmentation alone.
Our method also differs fundamentally from prior work such as SimTriplet, which uses spatial adjacency in histopathology but with a different formulation. SimTriplet employs a triplet architecture and custom loss function to enforce both intra- and inter-sample similarity, requiring a specialized model structure and training setup. In contrast, our method is architecture-agnostic and compatible with a wide range of SSL losses. It integrates directly into joint embedding frameworks such as BYOL, Barlow Twins, VICReg, and DINOv2 by combining augmentation-based and spatially adjacent positive pairs within a unified loss formulation. This design removes the need for handcrafted triplets or predictors, making our approach lightweight, flexible, and easy to adopt.
5.2 Relevance to Pathology Foundation Models
Recent advances in digital pathology have introduced large-scale foundation models that aim to generalize across datasets, tissue types, and diagnostic tasks [21, 26, 28]. These models typically rely on SSL objectives applied at massive scale. While our method is not intended as a replacement for such models, it offers a biologically motivated enhancement to existing pretraining strategies by incorporating spatial coherence into positive pair selection.
By aligning SSL sampling with the inherent architecture of tissue, our approach complements the goals of foundation models, particularly in settings where local morphological consistency is a strong learning signal. Its modular design allows for straightforward integration into pretraining pipelines without requiring architectural changes or significant computational overhead. As foundation models continue to scale, integrating spatial context into their learning objectives may further improve their data efficiency, interpretability, and clinical relevance.
5.3 Limitations and Future Directions
While our method demonstrates consistent performance improvements across datasets and SSL frameworks, several limitations remain. First, all experiments were conducted at a single magnification level using fixed patch sizes. Tissue morphology can exhibit scale-dependent features, particularly in structures with hierarchical organization. Future work should explore how spatial context-driven sampling behaves across multiple resolutions and its sensitivity to patch size and scale.
Second, while we evaluate the method across several SSL frameworks, each framework was tested with a single backbone architecture (e.g., ResNet-18 for Barlow Twins, BYOL, and VICReg; ViT-Tiny for DINOv2). Further testing with deeper or alternative backbones (e.g., ResNet-50, ConvNeXt, Swin Transformer) would provide stronger evidence for the approach’s framework- and architecture-agnostic applicability.
Third, our sampling strategy uses a fixed Chebyshev distance to define spatial neighbors. Although we conduct empirical ablation across multiple distances, a more adaptive approach—one that adjusts sampling radius based on local tissue density, slide size, or training dynamics—could yield more informative and robust positive pairs, particularly in heterogeneous or sparse tissues.
Fourth, due to computational constraints, we excluded DINOv2 from the spatial proximity experiments. Given the strong performance of transformer-based architectures in SSL, it remains an open and valuable direction to investigate how DINOv2 responds to spatial context-driven pair sampling.
Finally, our evaluation focuses primarily on classification tasks using attention-based MIL and linear probing. The potential benefits of spatially informed representations for other downstream tasks (such as clustering, retrieval, or feature interpretability) remain unexplored. Future research should examine how spatial context impacts representation disentanglement, interpretability, and performance in multi-task or clinical applications.
6 Conclusion
We proposed a spatial context-driven positive pair sampling strategy to enhance self-supervised learning (SSL) for whole-slide image (WSI) analysis. By leveraging the natural spatial coherence of tissue morphology, our approach introduces biologically grounded supervision into existing SSL pipelines without modifying their core architectures or objectives. Experimental results demonstrate that incorporating spatially adjacent positive pairs improves downstream performance, particularly in slide-level classification under Multiple Instance Learning (MIL) and patch-level classification via linear probing. Ablation studies show that smaller neighbor distances, especially immediate adjacency, yield the most consistent gains, though the optimal threshold can vary by dataset. Importantly, relying exclusively on spatial context may degrade representation quality, highlighting the value of balancing contextual and augmentation-based signals. Our goal was not to surpass state-of-the-art benchmarks, but rather to introduce a modular, interpretable improvement to SSL that reflects domain-specific structure. The findings suggest that spatial context is a valuable signal for learning more diagnostically relevant features in computational pathology. This strategy is particularly well-suited for large-scale, annotation-limited settings and could complement emerging pathology foundation models. Future work should explore multi-scale integration, adaptive sampling mechanisms, broader tissue diversity, and evaluation across additional tasks such as retrieval, segmentation, and clinical interpretability. Overall, our method offers a scalable and generalizable enhancement to self-supervised learning in histopathology.